# On the analysis of a higher-order Lotka-Volterra model: an application of $\mathcal{S}$-tensors and the polynomial complementarity problem

### v1

@@ IV Cooperative Lotka-Volterra Model
where for the given focal species $i$, $Q_{i}=\{j,k\,|\,(j,k)\neq(i,i)\}$, $x_{i}$ denotes the density of species $i$; $r_{i}$ is the per-capita intrinsic rate of increase of the focal species; $a_{ij}$ is the first-order coefficient, denoting $j$’s additive influence on $i$, and $b_{ijk}$ is the second-order (higher-order) coefficient, denoting $j$ and $k$’s joint non-additive influence on $i$, or alternatively $j$’s influence on $i$ correlated with a co-occurring third species $k$. Since we consider here cooperative systems, which represent the symbiosis among several species, it is natural to assume that the inter-specific interaction is non-negative. We further assume that the intra-specific interaction is non-positive, which is interpreted as competition within the species. Thus, all parameters in (30) are non-negative. We recall that a graph $G(A)$ is strongly connected if its adjacency matrix $A$ is irreducible. Throughout the section, we assume $A=[a_{ij}]$ is irreducible such that $G(A)$ is strongly connected.
@@
To derive the equilibrium set and to see whether it is finite, one only needs to check whether the equation set of $L_{i}=0$ (which is a set of quadratic equations with multiple variables) for $i\leq n$ has a finite number of solutions. If we set all $a_{ij}=0$ for $i\neq j$ and $b_{ijk}=0$ for $j,k\in Q_{i}$, it is straightforward to check that the equilibrium set is finite. Also note that this parameter setting can only be used to check whether there is a particular choice such that the equation has a finite number of solutions. Since this algebraic question is different from the analysis of a system, it doesn’t break the assumption that $A$ is irreducible. According to Theorem B.1 and Corollary B.2 in [22], since there exists a particular choice of parameters such that the equation has a finite number of solutions, if the parameters are generic and do not lie on a certain algebraic set of measure zero, then the equilibrium set is finite. ∎
Remark 8
@@
Throughout this paper, we say that the species is a winner when it takes some positive value in the corresponding equilibrium or is a loser when it takes the zero value. We use the set $S$ to denote the set of agents of the winner faction. A boundary equilibrium $X^{*}$ is an equilibrium where $x_{i}\neq 0$ for some $i\in S$ with non-empty $S$ and $x_{j}=0$ for the rest.
Theorem 9
@@
Recalling (30) and plugging $r_{i}x^{*}_{i}L_{i}(x^{*})=0$ in the equation above, one can get $(\mathbf{J}_{(X^{*})}X^{*})_{i}=-r_{i}x^{*}_{i}+r_{i}x^{*}_{i}(-b_{iii}x_{i}^{*2}+\sum_{j,k\in Q_{i}}b_{ijk}x^{*}_{j}x^{*}_{k})$. It follows that once there is equilibrium point $X^{*}=(x^{*}_{1},x^{*}_{2},\cdots,x^{*}_{n})\in\mathcal{T}$, and if $b_{ijk}$ ($i,j\in Q_{i}$) is sufficiently small, then $(\mathbf{J}_{(X^{*})}X^{*})_{i}<0$. Thus, the Jacobian is Hurwitz and furthermore, the equilibrium point $X^{*}=(x^{*}_{1},x^{*}_{2},\cdots,x^{*}_{n})$ is stable by Lemma 10.
Remark 9
Now, we know from Corollary 3 that the equilibrium corresponds to the solution of the polynomial complementarity problem. In light of Theorem 3, we can adjust the parameters in the following way. We decrease $b_{ijk}$ but also adjust $b_{iii}$ such that $\delta(-B)_{k}=-b_{kkk}-r_{k}(-B)_{-}$ remain unchanged. In this way, the upper bound of the equilibrium will be also unchanged. By continuously decreasing $b_{ijk}$, there must be a stable positive equilibrium. This technique is further applicable to the case of Theorem 13 below.
@@
System (30) has at least one positive equilibrium if $-B$ is an irreducible $\mathcal{M}$-tensor, $-A$ is an irreducible $\mathcal{M}$-matrix, and furthermore they are associated with the same positive vector $v$, i.e., $-Av>\mathbf{0}$ and $-Bv^{2}>\mathbf{0}$. If $A,B$ are furthermore both strictly diagonally dominant Metzler tensors with negative diagonal entries, then it is globally asymptotically stable.
Proof:
@@
Recall that an irreducible $\mathcal{M}$-tensor is a $\mathcal{S}$-tensor. ∎
@@
In Theorem 15.1.1 [33] and Theorem 14 [4], a sufficient condition of global stability of the unique positive equilibrium for a classical cooperative Lotka-Volterra model is provided. With the introduction of HOIs, positive equilibria of higher-order models are generally not unique. According to the definition of an irreducible monotone system, if there exists a positive equilibrium $X^{*}$, then $\{x|\mathbf{0}<x<X^{*}\}$ is positively invariant. This is a case when the system has an upper bound. Then, from Theorem 8, we know that for almost all initial conditions within $\{x|\mathbf{0}<x<X^{*}\}$, the solution converges to a positive equilibrium (not necessarily $X^{*}$, could be another, i.e. multi-stability may occur).
V Higher-order Two-faction-competition Lotka-Volterra Model
@@
The system of competition between two factions is also known as a system with limited competition. The system consists of two competing factions. The system generally follows a simple regulation, “friends’ friends are friends,” and “friends’ enemies are enemies”. Species within one faction cooperate with each other, while they compete with the species in the other faction.
@@
Proof:
Firstly, one can check that the Jacobian of (33)-(34) is of the form $\mathbf{J}=\left(\begin{array}[]{cc}M_{1}&T_{1}\\ T_{2}&M_{2}\\ \end{array}\right),$ where $M_{1}$, $M_{2}$ are irreducible Metzler matrices, and $T_{1}$, $T_{2}$ are non-positive matrices (it is irreducible as long as $z>\mathbf{0}_{n+m}$). It follows that $\mathbf{J}$ can be permuted into an irreducible Metzler matrix $\mathbf{\tilde{J}}=\left(\begin{array}[]{cc}M_{1}&-T_{1}\\ -T_{2}&M_{2}\\ \end{array}\right)$ via the matrix $P=\left(\begin{array}[]{cc}I_{m}&\mathbf{0}_{m\times n}\\ \mathbf{0}_{n\times n}&-I_{n}\\ \end{array}\right)$, so the system (33)-(34) is irreducible monotone. ∎
Remark 10
Theorem 10 tells us, in particular, that the Jacobian of the two-faction competition model can be permuted into an irreducible Metzler matrix. However, if one would deal with competition among more than two factions, the permutation may be no longer possible. For example, if we consider 3 factions, the structure of the Jacobian with 3 factions is analog to the case of Theorem 1 in [23]. A permutation is not possible for the same reason in [23]. In addition, the two-faction system may still diverge to infinity because of the cooperation terms.
In the following Theorems 11-15, we list all the possible equilibria of the model and study their stability.
Theorem 11
@@
Proof:
See Theorem 4. ∎
Theorem 12
@@
Proof:
Without loss of generality, we first investigate the case when the first faction is the winner. The equilibrium is then $(x^{*}_{1},x^{*}_{2},\cdots,x^{*}_{m},0,0,\cdots,0)=(X^{*},\mathbf{0}_{n})$. By plugging the equilibrium into the Jacobian, we obtain that $\mathbf{J}_{(X^{*},\mathbf{0}_{n})}=\left(\begin{array}[]{cc}M&\Omega\\ \mathbf{0}_{n\times m}&D\\ \end{array}\right),$ where $M$ is an irreducible Metzler matrix and represents the Jacobian of the sub-cooperative-system from the winner faction on an all-species-coexistence equilibrium point, $D$ is a diagonal matrix, and its diagonal entry reads $D_{i}=\hat{r}_{i}(1-\sum_{j\in S}\hat{b}_{ij}x_{j}^{*}-\sum_{j,k\in S}\hat{d}_{ijk}x_{j}^{*}x_{k}^{*}),\qquad i\in\mathbb{I}_{n}.$ Since the Jacobian is an upper-triangular block matrix, we know that the Jacobian is Hurwitz when all $D_{i}<0,i\in\mathbb{I}_{n}$ and the matrix $M$ is Hurwitz, which further implies that the coefficients of the first- and second-order competitive terms from the loser faction are sufficiently large and $(x^{*}_{1},x^{*}_{2},\cdots,x^{*}_{m})$ is a stable all-species-coexistence equilibrium point of the sub-cooperative-system from the winner faction when ignoring the loser faction. We recall that the second condition is satisfied when the cooperative HOIs terms are sufficiently small for the winners. The proof, for the case when the second faction is the winner, is exactly the same and thus omitted here. ∎
@@
Lemma 11 (Corollary 3.2 and Proposition 3.5 [24])
@@
Corollary 9
Consider the system (33)-(34), if $-A$ is an $\mathcal{H}^{+}$ matrix, $-B$ is an $\mathcal{H}^{+}$ tensor, and they are associated with the same positive vector $v$, i.e., $-Av>\mathbf{0}$ and $-Bv^{2}>\mathbf{0}$, then there exist at least one positive equilibrium. If $A,B$ are furthermore both strictly diagonally dominant tensors with negative diagonal entries, then it is globally asymptotically stable.
Proof:
Just to notice that a $\mathcal{H}^{+}$ tensor is an $\mathcal{S}$-tensor. ∎
Remark 11
Here, we are able to provide a strategy different from [5, 24] to obtain sufficient conditions for the global stability of the positive equilibrium.
Notice that $-A$ and $-B$ all have non-negative diagonal elements. The definition of an $\mathcal{H}^{+}$ tensor requires all positive diagonal elements. So, both setting are compatible with each other.
@@
According to Lemma 10, $\mathbf{\tilde{J}}$ is Hurwitz if $b_{ij},c_{ijk},d_{ijk},\hat{b}_{ij},\hat{c}_{ijk},\hat{d}_{ijk}$ are sufficiently small such that $\left(\mathbf{\tilde{J}}_{(X^{*},Y^{*})}\left(\begin{array}[]{c}X^{*}\\ Y^{*}\\ \end{array}\right)\right)_{i}<0$ for all $i$. ∎
Remark 12
In [5] (Theorem 3 and (A8) in Appendix) and [24] (Theorem 3.8), a sufficient condition for the global stability of a positive equilibrium (all-species-coexistence equilibrium) for the abstract system $\dot{N}_{i}=N_{i}F_{i}\left(N_{1},N_{2},\ldots,N_{m}\right),\quad i=1,2,\ldots,m$ was provided by Lyapunov function [5] or monotone system theory [24]. Since both systems proposed in our paper can be written in such a form, the results in [5] are also valid for our models. However, Theorem 3.8 [24] can not apply to our system because HOIs break the condition 3.2. The results in [5, 24] miss the possibility of bistability and what kind of role the HOIs play in the species’ coexistence. Our paper fills this gap. Moreover, the arguments similar to remark 9 is applicable for the case of Theorem 13.
Theorem 14
@@
Proof:
We firstly investigate the first case $(x_{1}^{*},\ldots,x_{l}^{*},\mathbf{0}_{m-l},\mathbf{0}_{n}),l<m$. The corresponding Jacobian matrix is $\mathbf{J}_{(X^{*},0,\cdots,0)}=\left(\begin{array}[]{cc}M&\Omega\\ \mathbf{0}_{(m+n-l)\times l}&D\\ \end{array}\right)$, where $M$ is an irreducible Metzler matrix and represents the Jacobian of the sub-cooperative-system from the winner faction on a boundary equilibrium point, and $D$ is a diagonal matrix. We know from Theorem 8, that $M$ is unstable and thus the equilibrium $(x_{1}^{*},\ldots,x_{l}^{*},\mathbf{0}_{m-l},\mathbf{0}_{n}),l<m$ is unstable. The proof, for the second case, is exactly analogous. ∎
Remark 13
@@
if $i$ denotes the loser agent in the first faction. Otherwise, if $i$ denotes the loser agent in the second faction, $D_{i}$ takes a similar form. In order to have all negative eigenvalues, the first- and second-order cooperative terms of the losers must be sufficiently small, so that all $D_{i}<0$, and $(x_{1}^{*},\cdots,x_{a}^{*},y_{1}^{*},\cdots,y_{b}^{*})$ must be a stable all-species-coexistence equilibrium point of the sub-system of the winners when ignoring the loser agents so that $M$ is Hurwitz. ∎
Remark 14
Theorem 15 further implies that any equilibrium, where the winners are in the different camps, but not all of them, is stable under the same condition as Theorem 15 because any of those equilibria can be permuted into the equilibrium $M^{*}$ or $N^{*}$ by index permutation. In light of Corollary 3, since the equilibrium will be bounded when $B$ has no-zero eigenvalue, we can confirm that there are always cases that conditions in Theorem 12,13, and 15 are satisfied.
VI Purely competitive higher-order Lotka-Volterra
In this section, we consider the case of competition among $n$ species: for $i\leq n$,

### v2

@@ 4 Cooperative Lotka-Volterra Model
where for the given focal species $i$, $Q_{i}=\{j,k\,|\,(j,k)\neq(i,i)\}$.
Since in this section we consider cooperative systems, which represent the symbiosis among several species, it is natural to assume that the inter-specific interaction is non-negative. We further assume that the intra-specific interaction is non-positive, which is interpreted as competition within the species. Thus, all parameters in (27) are non-negative. We recall that a graph $G(A)$ is strongly connected if its adjacency matrix $A$ is irreducible. Throughout the section, we assume $A=[a_{ij}]$ is irreducible such that $G(A)$ is strongly connected.
@@
To derive the equilibrium set and to see whether it is finite, one only needs to check whether the equation set of $L_{i}=0$ (which is a set of quadratic equations with multiple variables) for $i\leq n$ has a finite number of solutions. If we set all $a_{ij}=0$ for $i\neq j$ and $b_{ijk}=0$ for $j,k\in Q_{i}$, it is straightforward to check that the equilibrium set is finite. Also note that this parameter setting can only be used to check whether there is a particular choice such that the equation has a finite number of solutions. Since this algebraic question is different from the analysis of a system, it doesn’t break the assumption that $A$ is irreducible. According to [22, Theorem B.1 and Corollary B.2], since there exists a particular choice of parameters such that the equation has a finite number of solutions, if the parameters are generic and do not lie on a certain algebraic set of measure zero, then the equilibrium set is finite.
Remark 4.2.
@@
Throughout this paper, we say that a species is “a winner” when it takes some positive value in the corresponding equilibrium or is “a loser” when it takes the zero value. We use the set $S$ to denote the set of agents of the winner faction. A boundary equilibrium $X^{*}$ is an equilibrium where $x_{i}\neq 0$ for some $i\in S$ with non-empty $S$ and $x_{j}=0$ for the rest.
Theorem 14.
@@
Recalling (27) and plugging $r_{i}x^{*}_{i}L_{i}(x^{*})=0$ in the equation above, one can get $(\mathbf{J}_{(X^{*})}X^{*})_{i}=-r_{i}x^{*}_{i}+r_{i}x^{*}_{i}(-b_{iii}x_{i}^{*2}+\sum_{j,k\in Q_{i}}b_{ijk}x^{*}_{j}x^{*}_{k})$. It follows that once there is equilibrium point $X^{*}=(x^{*}_{1},x^{*}_{2},\cdots,x^{*}_{n})\in\mathcal{T}$, and if $b_{ijk}$ ($i,j\in Q_{i}$) is sufficiently small, then $(\mathbf{J}_{(X^{*})}X^{*})_{i}<0$. Thus, the Jacobian is Hurwitz and furthermore, the equilibrium point $X^{*}=(x^{*}_{1},x^{*}_{2},\cdots,x^{*}_{n})$ is stable by Lemma 4.3.
Remark 4.5.
Now, we know from Corollary 5 that the equilibrium corresponds to the solution of the polynomial complementarity problem. In light of Theorem 4, we can adjust the parameters in the following way to make sure that the condition of statement c) in Theorem 14 holds surely. We decrease $b_{ijk}$ but also adjust $b_{iii}$ such that $\delta(-B)_{k}=-b_{kkk}-r_{k}(-B)_{-}$ remain unchanged. In this way, the upper bound of the equilibrium will be also unchanged. By continuously decreasing $b_{ijk}$, there must be a stable positive equilibrium. This technique is further applicable to the case of Theorem 20 below.
@@
System (20) under the cooperative setting (i.e., (27)) has one unique positive equilibrium if $-B$ is an irreducible $\mathcal{M}$-tensor, $-A$ is an irreducible $\mathcal{M}$-matrix, and furthermore they are associated with the same positive vector $v$, i.e., $-Av>\mathbf{0}$ and $-Bv^{2}>\mathbf{0}$. From almost all initial conditions in $\{x|\mathbf{0}<x<x^{*}\}$, the solution converges to the unique positive equilibrium $x^{*}$. If $-Ax^{*}>\mathbf{0},-Bx^{*}>\mathbf{0}$ further holds, then the $x^{*}$ is globally asymptotically stable.
Proof 4.6.
@@
Recall that an irreducible $\mathcal{M}$-tensor is a $\mathcal{S}$-tensor. Thus, the system has a unique positive equilibrium according to Theorem 11.
Now, we prove the statement, from almost all initial conditions in $\{x|\mathbf{0}<x<x^{*}\}$, the solution converges to the unique positive equilibrium $x^{*}$. The cooperative system (27) is an irreducible monotone system. From the definition of irreducible monotone systems (section 2.3 [22]), we have $\phi_{t}(x)<\phi_{t}(x^{*})$ if $x<x^{*}$. Furthermore, if $x_{i}=0$, then $\dot{x}_{i}>0$. Combining both cases, we confirm that $\{x|\mathbf{0}<x<x^{*}\}$ is a positively invariant set of the system (27).
Then, we see that in $\{x|\mathbf{0}<x<x^{*}\}$, there are only one equilibrium, the $x^{*}$ on the boundary. According to Lemma 2.3 of [22], the solution converges to the unique positive equilibrium $x^{*}$ from almost all initial conditions in $\{x|\mathbf{0}<x<x^{*}\}$.
Next, we show the global stability when $-Ax^{*}>\mathbf{0},-Bx^{*}>\mathbf{0}$. Define the set $U_{1}=\{x|\max\frac{x_{i}}{x_{i}^{*}}\geq 1\}$ and $U_{2}=\{x|x\leq x^{*}\}$.
Then, we let $V_{m}=\max_{i}(\frac{x_{i}}{x^{*}_{i}})^{2}$ and $V=\max_{i}(\frac{x_{i}-x_{i}^{*}}{x^{*}_{i}})^{2}$. Further let $m=\arg\min_{i}(\frac{x_{i}}{x^{*}_{i}})^{2}$ and $t=\arg\min_{i}(\frac{x_{i}-x^{*}_{i}}{x^{*}_{i}})^{2}$. It holds $m=t$ because of $\frac{x_{i}-x^{*}_{i}}{x^{*}_{i}}=\frac{x_{i}}{x^{*}_{i}}-1$.
Furthermore, it must hold $x_{i}=\frac{x_{i}x_{i}^{*}}{x_{i}^{*}}\leq V_{m}^{\frac{1}{2}}x_{i}^{*}.$ It holds as a equality when $i=m$, otherwise it holds as an strict inequality.
Then, we have in the set $U_{1}$
$\begin{split}\dot{V}&=\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(Ax+Bx^{2}+\mathbf{1})_{m}\\ &=\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(\sum_{j}a_{mj}x_{j}+\sum_{jk}b_{mjk}x_{j}x_{k}+1)\\ &\leq\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(\sum_{j}a_{mj}x_{j}^{*}V_{m}^{\frac{1}{2}}+\sum_{jk}b_{mjk}x^{*}_{j}x^{*}_{k}V_{m}+1)\\ &\leq\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(\sum_{j}a_{mj}x_{j}^{*}V_{m}+\sum_{jk}b_{mjk}x^{*}_{j}x^{*}_{k}V_{m}+1)\\ &=\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(-V_{m}+1)\leq 0\quad\forall x\in U_{1}.\end{split}$
Otherwise, for the set $U_{2}$
$\begin{split}\dot{V}&\leq\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(\sum_{j}a_{mj}x_{j}^{*}V_{m}^{\frac{1}{2}}+\sum_{jk}b_{mjk}x^{*}_{j}x^{*}_{k}V^{\frac{1}{2}}_{m}+1)\\ &=\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(-V^{\frac{1}{2}}_{m}+1)\leq 0\quad\forall x\in U_{2}.\end{split}$
Notice that $V$ is locally positive definite in $U_{1}$ and $U_{2}$, and $\dot{V}$ is negative semidefinite. The solution must converge to $\{x|\dot{V}=0\}$. In order to have $\dot{V}=0$, for every $i$, it must hold $i=m$ and $x_{m}=x_{m}^{*}$ which indicates that $\{x|\dot{V}=0\}$ only contains $x^{*}$. Thus, $x^{*}$ is asymptotically stable with a domain of attraction $U_{1},U_{2}$.
Next, we further let $L_{s}=\min_{i}(\frac{x_{i}}{x^{*}_{i}})^{2}$ and $s=\arg\min_{i}(\frac{x_{i}}{x^{*}_{i}})^{2}$.
Furthermore, it must hold $x_{i}=\frac{x_{i}x_{i}^{*}}{x_{i}^{*}}\geq L_{s}^{\frac{1}{2}}x_{i}^{*}.$ It holds as a equality when $i=s$, otherwise it holds as an strict inequality.
Then, consider the set $U_{3}=\{x|\min_{i}\frac{x_{i}}{x^{*}_{i}}<1\}$, we have
$\begin{split}\dot{V}&=\frac{2r_{s}x_{s}}{(x_{s}^{*})^{2}}(x_{s})(Ax+Bx^{2}+\mathbf{1})_{s}\\ &=\frac{2r_{s}x_{s}^{2}}{(x_{s}^{*})^{2}}(\sum_{j}a_{sj}x_{j}+\sum_{jk}b_{sjk}x_{j}x_{k}+1)\\ &\geq\frac{2r_{s}x_{s}^{2}}{(x_{s}^{*})^{2}}(\sum_{j}a_{sj}x_{j}^{*}L_{s}^{\frac{1}{2}}+\sum_{jk}b_{sjk}x^{*}_{j}x^{*}_{k}L_{s}+1)\\ &\geq\frac{2r_{s}x_{s}^{2}}{(x_{s}^{*})^{2}}(\sum_{j}a_{sj}x_{j}^{*}L_{s}+\sum_{jk}b_{sjk}x^{*}_{j}x^{*}_{k}L_{s}+1)\\ &=\frac{2r_{s}x_{s}^{2}}{(x_{s}^{*})^{2}}(-L_{s}+1)>0\quad\forall x\in U_{3}.\end{split}$
This guarantees that all trajectories in $U_{3}$ finally enter $U_{1}$. Thus, the $x^{*}$ is globally asymptotically stable.
@@
In [35, Theorem 15.1.1] and [4, Theorem 14], a sufficient condition of global stability of the unique positive equilibrium for a classical cooperative Lotka-Volterra model is provided. With the introduction of HOIs, positive equilibria of higher-order models are generally not unique. According to the definition of an irreducible monotone system, if there exists a positive equilibrium $X^{*}$, then $\{x|\mathbf{0}<x<X^{*}\}$ is positively invariant. This is a particular case when the system has an upper bound. Then, from Theorem 13, we know that for almost all initial conditions within $\{x|\mathbf{0}<x<X^{*}\}$, the solution converges to a positive equilibrium (not necessarily $X^{*}$, could be another, i.e. multi-stability may occur).
5 Higher-order Two-faction-competition Lotka-Volterra Model
@@
A system of competition between two factions is also known as a system with limited competition [24]. The system generally follows a simple regulation, “friends’ friends are friends,” and “friends’ enemies are enemies”. Species within one faction cooperate with each other, while they compete with the species in the other faction.
@@
Proof 5.1.
Firstly, one can check that the Jacobian of (30)-(31) is of the form $\mathbf{J}_{z}=\left(\begin{array}[]{cc}M_{1}&T_{1}\\ T_{2}&M_{2}\\ \end{array}\right),$ where $M_{1}$, $M_{2}$ are irreducible Metzler matrices, and $T_{1}$, $T_{2}$ are non-positive matrices (it is irreducible as long as $z>\mathbf{0}_{n+m}$). It follows that $\mathbf{J}$ can be permuted into an irreducible Metzler matrix $\mathbf{\tilde{J}}=\left(\begin{array}[]{cc}M_{1}&-T_{1}\\ -T_{2}&M_{2}\\ \end{array}\right)$ via the matrix $P=\left(\begin{array}[]{cc}I_{m}&\mathbf{0}_{m\times n}\\ \mathbf{0}_{n\times n}&-I_{n}\\ \end{array}\right)$, so the system (30)-(31) is irreducible monotone.
Remark 5.2.
Theorem 16 tells us, in particular, that the Jacobian of the two-faction competition model can be permuted into an irreducible Metzler matrix. However, if one would deal with competition among more than two factions, the permutation may be no longer possible. For example, if we consider 3 factions, the structure of the Jacobian with 3 factions is analog to the case of [23, Theorem 1]. A permutation is not possible for the same reason in [23]. In addition, the two-faction system may still diverge to infinity because of the cooperation terms.
In the following Theorems 17-22, we list all the possible equilibria of the model and study their stability.
Theorem 17.
@@
Proof 5.3.
See Theorem 6.
Theorem 18.
@@
Proof 5.4.
Without loss of generality, we first investigate the case when the first faction is the winner. The equilibrium is then $(x^{*}_{1},x^{*}_{2},\cdots,x^{*}_{m},0,0,\cdots,0)=(X^{*},\mathbf{0}_{n})$. By plugging the equilibrium into the Jacobian, we obtain that $\mathbf{J}_{(X^{*},\mathbf{0}_{n})}=\left(\begin{array}[]{cc}M&\Omega\\ \mathbf{0}_{n\times m}&D\\ \end{array}\right),$ where $M$ is an irreducible Metzler matrix and represents the Jacobian of the sub-cooperative-system from the winner faction on an all-species-coexistence equilibrium point, $D$ is a diagonal matrix, and its diagonal entry reads $D_{i}=\hat{r}_{i}(1-\sum_{j\in S}\hat{b}_{ij}x_{j}^{*}-\sum_{j,k\in S}\hat{d}_{ijk}x_{j}^{*}x_{k}^{*}),\qquad i\in\mathbb{I}_{n}.$ Since the Jacobian is an upper-triangular block matrix, we know that the Jacobian is Hurwitz when all $D_{i}<0,i\in\mathbb{I}_{n}$ and the matrix $M$ is Hurwitz, which further implies that the coefficients of the first- and second-order competitive terms from the loser faction are sufficiently large and $(x^{*}_{1},x^{*}_{2},\cdots,x^{*}_{m})$ is a stable all-species-coexistence equilibrium point of the sub-cooperative-system from the winner faction when ignoring the loser faction. We recall that the second condition is satisfied when the cooperative HOIs terms are sufficiently small for the winners. The proof, for the case when the second faction is the winner, is exactly the same and thus omitted here.
@@
Lemma 5.5 (Corollary 3.2 and Proposition 3.5 [24]).
@@
Corollary 19.
Consider the system (30)-(31), if $-A$ is an $\mathcal{H}^{+}$ matrix, $-B$ is an $\mathcal{H}^{+}$ tensor, and they are associated with the same positive vector $v$, i.e., $-Av>\mathbf{0}$ and $-Bv^{2}>\mathbf{0}$, then there exist one unique positive equilibrium.
Proof 5.6.
It suffices to notice that a $\mathcal{H}^{+}$ tensor is an $\mathcal{S}$-tensor.
Notice that $-A$ and $-B$ have non-negative diagonal elements. The definition of an $\mathcal{H}^{+}$ tensor requires all positive diagonal elements. So, both settings are compatible with each other. Furthermore, once the negative diagonal entries of $a_{ii},b_{iii}$ are sufficiently large (in absolute value) compared with the off-diagonals, then the system has a unique globally asymptotically stable equilibrium from Theorem 11 and Corollary 10.
@@
According to Lemma 4.3, $\mathbf{\tilde{J}}$ is Hurwitz if $b_{ij},c_{ijk},d_{ijk},\hat{b}_{ij},\hat{c}_{ijk},\hat{d}_{ijk}$ are sufficiently small such that $\left(\mathbf{\tilde{J}}_{(X^{*},Y^{*})}\left(\begin{array}[]{c}X^{*}\\ Y^{*}\\ \end{array}\right)\right)_{i}<0$ for all $i$.
Remark 5.8.
In [5] (Theorem 3 and (A8) in Appendix) and [24] (Theorem 3.8), a sufficient condition for the global stability of a positive equilibrium (all-species-coexistence equilibrium) for the abstract system $\dot{N}_{i}=N_{i}F_{i}\left(N_{1},N_{2},\ldots,N_{m}\right),\quad i=1,2,\ldots,m$ was provided by Lyapunov function [5] or monotone system theory [24]. Since both systems proposed in our paper can be written in such a form, the results in [5] are also valid for our models. However, Theorem 3.8 [24] can not be applied to our system because HOIs break condition 3.2 in [24]. The results in [5, 24] miss the possibility of bistability and what kind of role the HOIs play in the species’ coexistence. Our paper fills this gap. Moreover, the arguments similar to remark 4.5 are applicable for the case of Theorem 20. For the global stability, the corollary 10 is directly applicable.
Theorem 21.
@@
Proof 5.9.
We firstly investigate the first case $(x_{1}^{*},\ldots,x_{l}^{*},\mathbf{0}_{m-l},\mathbf{0}_{n}),l<m$. The corresponding Jacobian matrix is $\mathbf{J}_{(X^{*},0,\cdots,0)}=\left(\begin{array}[]{cc}M&\Omega\\ \mathbf{0}_{(m+n-l)\times l}&D\\ \end{array}\right)$, where $M$ is an irreducible Metzler matrix and represents the Jacobian of the sub-cooperative-system from the winner faction on a boundary equilibrium point, and $D$ is a diagonal matrix. We know from Theorem 13, that $M$ is unstable and thus the equilibrium $(x_{1}^{*},\ldots,x_{l}^{*},\mathbf{0}_{m-l},\mathbf{0}_{n}),l<m$ is unstable. The proof, for the second case, is exactly analogous.
Remark 5.10.
@@
if $i$ denotes the loser agent in the first faction. Otherwise, if $i$ denotes the loser agent in the second faction, $D_{i}$ takes a similar form. In order to have all negative eigenvalues, the first- and second-order cooperative terms of the losers must be sufficiently small, so that all $D_{i}<0$, and $(x_{1}^{*},\cdots,x_{a}^{*},y_{1}^{*},\cdots,y_{b}^{*})$ must be a stable all-species-coexistence equilibrium point of the sub-system of the winners when ignoring the loser agents so that $M$ is Hurwitz.
Remark 5.12.
Theorem 22 further implies that any equilibrium, where the winners are in the different camps, but not all of them, is stable under the same condition as Theorem 22 because any of those equilibria can be permuted into the equilibrium $M^{*}$ or $N^{*}$ by index permutation. In light of Corollary 5 and Lemma 2.9, by following a similar procedure of Remark 4.5, we can confirm that there are always cases that conditions in Theorem 18, 20, and 22 are satisfied.
6 Purely competitive higher-order Lotka-Volterra
In this section, we consider the case of competition among $n$ species: